OpenStudy (anonymous):

Solve by substitution. x+7y=45 2x+7y=41

OpenStudy (zarkon):

from the first equation x=45-7y put this into the second equation 2(45-7y)+7y=41 solve for y plug value of y back into x=45-7y to get value of x

OpenStudy (zarkon):

you should get x=-4 and y=7

OpenStudy (agentjamesbond007):

Take the first equation and subtract the 7y to both sides and gives you x=-7y+45 next take the other equation and plug it in to make 2(-7y+45)+7y=41 distribute -14y+90+7y=41 combine like terms -7y+90=41 subtract 90 to both sides -7y=49 divide by -7 to both sides y=-7 that is your y-coordinate next, plug it into the other equation to get the x

OpenStudy (zarkon):

"subtract 90 to both sides -7y=49" -7y=-49